OpenStudy (anonymous):

MEDAL :) Please help me Which ratios do not form a proportion? A. 24/96 , 27/108 B. 17/102 , 12/70 C. 21/51 , 35/85 B. 45/75 , 63/105

OpenStudy (anonymous):

@I_Need_Help_With_Home

OpenStudy (anonymous):

please help me @aum

OpenStudy (aum):

If a/b = c/d then, a*d = b*c (cross-multiplying). So IF 24/96 = 27/108 then check if 24*108 is equal to 27*96

OpenStudy (anonymous):

I got 2916

OpenStudy (anonymous):

I did 108 x 24

OpenStudy (aum):

If 24 x 108 = 27 x 96 then 24/96 and 27/108 form a proportion.

OpenStudy (anonymous):

But I need to know which ratios do not form a proportion

OpenStudy (anonymous):

Ok then so I multiply the same numbers with each other choices

OpenStudy (aum):

Correct. Try the same method with each answer choice and see which one is not true.

OpenStudy (anonymous):

So would it be C.

OpenStudy (anonymous):

@aum

OpenStudy (aum):

B. 17/102 , 12/70 17 x 70 = 1190 12 x 102 = 1224 1190 is not equal to 1224. Therefore 17/102 and 12/70 do NOT form a proportion. Try the other two choices too just to make sure.

OpenStudy (aum):

All you have to do is cross-multiply and check if they are the same.

OpenStudy (aum):

\[ \frac{21}{51}, ~~~\frac{35}{85} \\ \text{Cross-multiply:} \\ 21 \times 85 = 1785 \\ 35 \times 51 =1785 \\ \text{Both are the same. Therefore, they form a proportion.} \]